# Temporal Test-Time Adaptation with State-Space Models

### v2

@@ D.3 Cluster Visualization
D.4 Runtime
Table 9: Relative runtime per batch compared to the source model: BN and T3A are the fastest for adaptation, while STAD takes significantly less time than CoTTA, which updates all model parameters. Note that baselines like TENT and CoTTA benefit from highly optimized backpropagation code, whereas our codebase has not been optimized for speed. Improving the efficiency of our code would make STAD likely faster.
Methods Yearbook FMoW-Time
Source Model 1.0 1.0
Batch Norm (BN) 1.0 1.1
TENT 1.4 6.4
CoTTA 17.1 200
SHOT 1.3 6.3
SAR 1.5 7.1
T3A 1.1 1.8
LAME 1.2 2.9
STAD-vMF 2.5 30.8
STAD-Gauss 3.3 -
D.5 Sensitivity to Hyperparameters
Table 10: Sensitivity of concentration parameters $\kappa^{\mathrm{trans}}$ and $\kappa^{\mathrm{ems}}$ for STAD-vMF on Yearbook (uniform label distribution) in accuracy: In general, high concentration parameters model more compact clusters ($\kappa^{\mathrm{ems}}$) and limited prototype movement ($\kappa^{\mathrm{trans}}$). Low concentration parameters allow for loose clusters and faster dynamics.
Value $\kappa^{\mathrm{trans}}$ $\kappa^{\mathrm{ems}}$
50 85.5034 $\pm$ 1.3031 85.5022 $\pm$ 1.3036
100 85.5034 $\pm$ 1.3031 85.5034 $\pm$ 1.3031
500 85.5034 $\pm$ 1.3031 85.5043 $\pm$ 1.3019
1000 85.5034 $\pm$ 1.3031 85.5058 $\pm$ 1.3001
5000 85.4980 $\pm$ 1.3099 85.5058 $\pm$ 1.3001
Table 11: Sensitivity of sliding window parameter $s$ for STAD-vMF on Yearbook (uniform label distribution): A smaller $s$ reduces the influence of past prototypes, while a larger $s$ extends the considered history, giving more weight to past prototypes. However, this comes at the cost of increased computational burden, as runtime scales linearly with window size. In our experiments, we use a default window size of $s=3$. The results suggest that giving more weight to past inferred prototypes may further improve performance.
$s$ Accuracy
3 85.4975 $\pm$ 1.34
5 85.5022 $\pm$ 1.30
7 85.5029 $\pm$ 1.31

### v3

@@ E.6 Cluster Visualization
E.7 Computational Complexity and Runtime
The complexity of STAD-vMF scales linearly with the sliding window size $s$, batch size $N_{t}$, representation dimension $D$ and number of classes $K$, i.e. $\mathcal{O}(s\times N_{t}\times D\times K)$. Note that the sliding window size $s$ is fixed at $s=3$ and operations over the batch size, number of classes and representation dimension can be fully parallelized.
Tab.˜13 reports relative runtime compared to the source model for STAD and baselines. As results on ImageNet-V2 show, even for high dimensions ($D=2048$) and large number of classes ($K=1000$) STAD’s runtime is comparable to baseline TTA methods.
Table 13: Relative runtime per batch compared to the source model
Methods Yearbook FMoW-Time ImageNetV2
Source Model 1.0 1.0 1.0
BN 1.0 1.1 1.1
TENT 1.4 6.4 7.2
CoTTA 17.1 200.0 327.3
SHOT 1.3 6.3 6.7
SAR 1.5 7.1 10.0
RoTTA 90.0 30.0 41.8
LAME 1.2 2.9 5.5
T3A 1.1 1.8 10.9
STAD-vMF 2.5 8.8 23.6
STAD-Gauss 3.3 - -
E.8 Sensitivity to Hyperparameters
In this section, we conduct a sensitivity analysis of the hyperparameters involved in STAD. We analyze sensitivity on two datasets: the temporal shift dataset Yearbook and the commonly used corruption benchmark CIFAR-10-C. All experiments are conducted under a uniform label distribution. When testing the sensitivity to a specific hyperparameter, all other hyperparameters are fixed at their default values (see App.˜D). Results show the average over three random training seeds for Yearbook and the average over 15 corruption types for CIFAR-10-C.
Sensitivity to sliding window size $s$
The window size determines the number of past time steps considered by the dynamic model. A small $s$ limits the influence of past prototypes, whereas a larger $s$ extends the considered history, giving more weight to past prototypes. However, large values of $s$ come at the cost of increased computational burden, as runtime scales linearly with window size. Tab.˜14 suggests that increasing the window size could improve adaptation performance.
Table 14: Accuracy of STAD for different values of $s$
$s$ 3 5 7
Yearbook 85.4975 $\pm$ 1.34 85.5022 $\pm$ 1.30 85.5029 $\pm$ 1.31
CIFAR-10-C 76.9683 $\pm$ 11.25 76.9735 $\pm$ 11.25 76.9823 $\pm$ 11.24
Sensitivity to $\kappa^{trans}$
The transition concentration parameter $\kappa^{trans}$ regulates the transition noise and determines how far cluster prototypes move between different time steps. A high concentration value $\kappa^{trans}$ implies little movement of class prototypes, whereas low $\kappa^{trans}$ allows prototypes to move more. This parameter thus acts as a regularization factor between a more static and a more dynamic model. Tab.˜15 displays the results. Performance changes only marginally for different values of the concentration parameter.
Table 15: Accuracy of STAD-vMF for different values of $\kappa^{\mathrm{trans}}$
$\kappa^{\mathrm{trans}}$ 50 100 500 1000 5000
Yearbook 85.5034 $\pm$ 1.3031 85.5034 $\pm$ 1.3031 85.5034 $\pm$ 1.3031 85.5034 $\pm$ 1.3031 85.4980 $\pm$ 1.3099
CIFAR-10-C 76.9684 $\pm$ 11.2540 76.9683 $\pm$ 11.2543 76.9685 $\pm$ 11.2538 76.9685 $\pm$ 11.2538 76.9688 $\pm$ 11.2552
Sensitivity to $\kappa^{ems}$
The emission concentration parameter $\kappa^{ems}$ regulates the emission noise and determines the spread of clusters. A high concentration value $\kappa^{ems}$ implies small, compact clusters, while low $\kappa^{ems}$ allows for widespread clusters. Results are shown in Tab.˜16.
Table 16: Accuracy of STAD-vMF for different values of $\kappa^{\mathrm{ems}}$
$\kappa^{\mathrm{ems}}$ 50 100 500 1000 5000
Yearbook 85.5022 $\pm$ 1.3036 85.5034 $\pm$ 1.3031 85.5043 $\pm$ 1.3019 85.5058 $\pm$ 1.3001 85.5058 $\pm$ 1.3001
CIFAR-10-C 76.9679 $\pm$ 11.2543 76.9683 $\pm$ 11.2543 76.9689 $\pm$ 11.2529 76.9700 $\pm$ 11.2514 76.9700 $\pm$ 11.2513